OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log 1/5 x. Part 1: Describe the domain, range, and general shape of this graph. Part 2: Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts.

OpenStudy (anonymous):

pick points that are easy to plot

OpenStudy (anonymous):

for example if \(x=1\) then \(\log_{1/5}(1)=0\)

OpenStudy (anonymous):

then try powers of 5, like for example \[\log_{1/5}(5)=-1\] since \((\frac{1}{5})^{-1}=5\)

OpenStudy (ilovebmth1234):

The log function has vertical asymptote at x=0 x-intercept at x=1 Because its base 1/5, when x =1/5, y = 1 that should give the general shape of the graph

OpenStudy (anonymous):

if you want a general picture, click here http://www.wolframalpha.com/input/?i=log_ {1%2F5}%28x%29 make sure to choose "real valued plot" so you don't get messed up with complex numbers the domain of any log function is \(x>0\) and the range is \((-\infty, \infty)\)

OpenStudy (mathmale):

It's late in the game to be pointing this out, but your f(x) = log 1/5 x. is subject to varying interpretations. Do you mean \[\log_{1/5}x~or~\log \frac{ 1 }{ 5 }x ~or~what? \]